OpenStudy (anonymous):

find the slope-intercept equation to find the line that given characteristics slope 9.4and y-intercept (0,-5)

OpenStudy (anonymous):

You mean you want to find the equation right? If that is the question then you can plug in the given you have in the following general equation:\[y-yo= m(x-xo)\] and you'll get: \[y +5 = 9.4(x-0) \rightarrow y =9.4x -5\] Hope everything is clear now ^_^